# Hybrid plasmonic nanoantennas using corner reflectors for enhanced single-photon emission

### v1

@@ 3 Results and Discussion
One of the main differences observed at small film thickness ($W=10$ nm) is the presence of relevant amplitude side lobes at the lower values of $S$. This phenomenon leads to reduced directivity values, despite the main lobe exhibiting an narrower beamwidth at $W=10$ nm. Conversely, the patterns seen at $W=40$ nm and $W=80$ nm are qualitatively similar, indicating a convergence as the thickness $W$ increases.
For thicker sheet CRs, the gradual narrowing (or broadening) of the on-axis lobe contributes to regions of increasing (or decreasing) directivity values. The zero observed in the directivity curves arises from the splitting of the main lobe at $S=\lambda_{M}$ that produces a radiation null along the $Y$ axis. In contrast, for $W=10$, the local minima in the directivity plots correlate with the increasing relative weight of the side lobes in the radiation pattern. As $S$ increases, separation of the main lobe is also noted; however, it occurs along the horizontal plane, unlike the vertical splitting seen at $W=40$ nm and $W=80$ nm. As $S$ continues to grow, additional lobe bifurcations become apparent, leading to the emergence of a fractal fractal structure in the radiation pattern for very large values of $S$ in extensive CRs, as analyzed in [18].
The sequences of radiation diagrams presented in Figs. 6, 7, and 8 are very similar for the corresponding ones of gold CRs when the values of $S$ are normalized by the respective operation wavelength $\lambda_{M}$.
We now examine the impact of the corner reflector on the losses and the emission rate of the dipole-NP system. Fig. 9 displays $f_{R}$, and the normalized loss factor $f_{L}$ of NAs incorporating silver CRs of four different thicknesses, $W$, as a function of the normalized distance, $S/\lambda_{Ag}$, between the silver NP and the corner apex. The corresponding results for gold NAs are shown in Fig. 10. In all the figures, the horizontal lines pinpoint the reference values for the NPs without the CRs.
@@
Figure 9: Radiative transition rate enhancement $f_{R}$ and normalized loss $f_{L}$ factors of a silver CR as a function of the normalized feeder position $S/\lambda_{M}$ for four different reflector film thicknesses: (a) $W=10$ nm, (b) $W=20$ nm, (c) $W=40$ nm and (d) $W=80$ nm. The horizontal lines describe the corresponding reference levels without the CR.
@@
Figure 10: Radiative transition rate enhancement $f_{R}$ and normalized loss $f_{L}$ factors of a gold CR as a function of the feeder position $S/\lambda_{M}$ for four different reflector film thicknesses: (a) $W=10$ nm, (b) $W=20$ nm, (c) $W=40$ nm and (d) $W=80$ nm.
@@
For silver CRs, the normalized loss factors display moderate deviations from the reference values except for very small values of $S/\lambda_{Ag}$, where a steep increase of the losses occurs due to stronger interactions with the nearby CR metal films when the thickness, $W$, is sufficiently large. In contrast, for gold CRs, Fig. 10 indicates larger fluctuations in the loss factor around the reference levels.
@@
The calculated radiative emission rate enhancement factors also exhibit a pronounced effect due to the interaction with the corner sheets when $W$ is sufficiently large. This interaction significantly impacts both silver and gold NAs, and the associated $S/\lambda_{M}$ range is larger than that observed for $f_{L}$ in silver NRs, increasing with $W$. The variation with $S/\lambda_{M}$ follows a distinct pattern: $f_{R}$ reaches a maximum near the first peak of directivity, then drops to a minimum around the directivity null at $S/\lambda_{M}=1$, before rising again near the second peak.
@@
For silver CRs with $W=80$ nm, we find that near the first directivity maximum at $S=0.30$ $\mu$m ($S/\lambda_{Ag}=0.67$), the values are $D=20.1$, $f_{R}=14.5$, and $f_{L}=6.7$. The second maximum yields a significantly higher directivity of $D=40.4$ at $S=0.55$ $\mu$m ($S/\lambda_{Ag}=1.22$), with $f_{R}=6.0$ and $f_{L}=5.9$. Thus, the CR provides very large improvements of the directivity when compared to that of a electric dipole emitter, $D=1.5$.
Considering the reference values for the NP emitter, $f_{R}=8.3$ and $f_{L}=5.4$, the first maximum offers a significant increase in both directivity and radiative emission rate enhancement factor, with only a modest rise in losses. At the second maximum, the directivity is much larger, though there is a slight reduction in $f_{R}$ in relation to the reference. The loss factor, however, remains close to that of the emitter without the CR. The gain figure of merit is $G=13.8$ and $G=20.4$ for the first and second maxima, respectively, largely exceeding the value $G=0.9$ obtained for the NA without the CR.
For the gold NA, the reference values $f_{R}=4.9$ and $f_{L}=6.1$. With a $W=80$ nm CR NA, at the first maximum, located at $S=0.35$ $\mu$m ($S/\lambda_{Au}$=0.58), the directivity is $D=18.0$, with $f_{R}=8.2$ and $f_{L}=7.0$. The second maximum occurs at $S=0.75$ $\mu$m ($S/\lambda_{Au}=0.58$), where $D=29.0$, $f_{R}=4.4$ and $f_{L}=6.7$.
For gold CRs, the first maximum provides a high directivity and an improvement of $f_{R}$. At the second maximum, while the directivity is even greater, $f_{R}$ is close to the reference value. In both cases, the losses are similar to those of the gold NP without the CR. The gain figure of merit is $G=8.0$ and $G=11.5$ for the first and second maxima, respectively, significantly exceeding the value $G=0.7$ obtained for the NA without the CR but significantly lower than the values obtained for silver CR NAs. Besides the impact of the better figures for $f_{R}$ and $f_{L}$ in the case of silver NAs, the lower values of $G$ obtained in the case of gold CR NAs is also due to the fact that $L/\lambda_{M}$ is smaller in the case of gold.
@@
Figure 11: Values obtained of the directivity as a function of the feeder position $S$ normalized to the corresponding target wavelength for three different reflector film thicknesses when the length of the CR planes is increased to $L=3$ $\mu$m. (a) Results for silver CR. (b) Results for gold CR.
@@
Fig. 11 shows the directivities calculated for silver and gold CR NAs when the corner length is increased to $L=3$ $\mu$m for three values of $W$ ($10$ nm, $40$ nm and $80$ nm). For this $L$, the peak directivities are essentially the same for silver and gold CRs, even though $L/\lambda_{M}$ is larger for silver. This indicates a saturating effect in the increase of $D$ with $L$.
@@
In the case of gold, the relative positions of the maxima and minima as a function of $S/\lambda_{M}$ closely match those found for $L=1.5$ $\mu$m, with a null occurring at $S=\lambda_{M}$ for $W=80$ nm. In contrast, for silver, significant shifts in the directivity maxima and minima as a function of $S/\lambda_{M}$ are observed , and these shifts are highly dependent on the value of $W$.
@@
Figure 12: Radiative transition rate enhancement $f_{R}$ and normalized loss $f_{L}$ factors of a gold CR with $L=3$ $\mu$m as a function of the feeder position $S/\lambda_{M}$ for three different reflector film thicknesses: (a) silver CR $W=10$ nm, (b) gold CR $W=10$ nm, (c) silver CR $W=40$ nm, (d) gold CR $W=40$ nm, (d) silver CR $W=80$ nm, (d) gold CR $W=80$ nm.
The normalized radiative emission and loss rates for $L=3$ $\mu$m CRs are displayed in Fig. 12 for the three values of $W$ considered. The results closely resemble those obtained for $L=1.5$ $\mu$m, indicating that the impact of the CR on the loss and emission properties is due to the interactions occurring in a localized region near the apex of the CR.
Fig. 13 shows the gain, $G$, of silver and gold CR NAs as a function of $S/\lambda_{M}$. The gain provides a comprehensive measure of on-axis photon emission, incorporating all three key parameters: $f_{R}$, $f_{L}$, and $D$. For gold CR NAs, the maximum gain occurs at a metal thickness of $W=80$ nm. The first gain peak, with a value of $G=11.8$, is found at $S/\lambda_{Au}=0.67$ and $S/\lambda_{Au}=0.75$. The second peak, $G=21.8$, is located at $S/\lambda_{M}=1.33$.
In the case of silver, the aforementioned shift of the directivity maxima with respect to $S/\lambda_{Ag}$ results in the highest directivity for $W=40$ nm rather than $W=80$ nm. This shift aligns the directivity peak with a higher $f_{R}$ value at $W=40$ nm, leading to the maximum gain being achieved at this thickness. The first gain peak for silver is $G=12.3$ at $S/\lambda_{Ag}=0.5$ and the second is $G=30.0$ at $S/\lambda_{Ag}=1.25$.
@@
Figure 13: Nanoantenna gain as a function of the feeder position $S$ normalized to the corresponding target wavelength for three different reflector film thicknesses when the length of the CR planes is increased to $L=3$ $\mu$m. (a) Results for silver CR. (b) Results for gold CR.
@@ 4 Conclusions
The analysis of dipole-NP interactions, combined with CR integration, reveals important insights into the enhancement of radiative properties for silver and gold NPs. The fundamental behavior of NP interactions is crucial for understanding the modifications introduced by the incorporation of the CR in the system. Initially, isolated dipole-NP calculations without the CR showed trends consistent with previous theoretical and experimental studies, particularly regarding the variation in resonance wavelength, the radiative emission rate enhancement factor, and loss factor with silver and gold particle size and dipole emitter separation. This preliminary study established a set of parameters for the subsequent analysis of the CR. The results also suggest that silver NPs are more effective than gold NPs in enhancing radiative emission while maintaining lower loss factors.
Integrating CRs into silver and gold NA system leads to substantial improvements in directivity. The variation in directivity with respect to the normalized distance $S/\lambda_{M}$ reveals that thicker metal films generally lead to more regular oscillatory patterns in directivity. At the first directivity maxima, significant gains in the radiative enhancement factor, $f_{R}$, are also achieved with minimal increases in the loss factor, $f_{L}$, leading to high gain figures of merit, $G$. At the second directivity maximum, the CR mainly enhances directivity, with the highest values of $D$ corresponding to the global maxima of $G$.
Extending the corner length to $L=3$ $\mu$m demonstrates a saturation effect in the enhancement of directivity for both materials, with similar peak values achieved for gold and silver CRs despite differences in $L/\lambda_{M}$. Meanwhile, the patterns for the variation of $f_{R}$ and $f_{L}$ closely resemble those found at $L=1.5$ $\mu$m. These findings highlight the localized nature of CR-induced enhancements near the apex. However, silver CRs exhibit larger shifts in the positions of directivity maxima and minima with changes in $W$. This results in the maximum gain for silver CR NAs at a metal thickness $W=40$ nm, rather than $W=80$ nm, due to the alignment of the directivity maximum with a higher value of $f_{R}$.
Overall, the results indicate that silver CR NAs outperform gold CR NAs in radiative enhancement and gain, offering more efficient light emission and directivity control for nanoscale optical applications, although different spectral windows are targeted by the choice of NA metal.
The reflector angle selected for the analysis aligns with a typical value used in RF applications, corresponding to optimization for operation at the more favorable second gain maximum. The analysis reveals a strong metal-feeder interaction near the corner apex, suggesting the potential benefits of further exploring corner geometries to optimize the first gain maximum, which will be addressed in future works.

### v2

@@ 3 Results and Discussion
Figure 9: Purcell, $f_{R}$, and normalized loss, $f_{L}$, factors of silver CR as a function of the normalized feeder position $S/\lambda_{M}$ for four different reflector film thicknesses: (a) $W=10$ nm, (b) $W=20$ nm, (c) $W=40$ nm, and (d) $W=80$ nm. The horizontal lines describe the corresponding reference levels without the CR.
@@
Figure 10: Radiative transition rate enhancement, $f_{R}$, and normalized loss, $f_{L}$, factors of gold CR as a function of the normalized feeder position $S/\lambda_{M}$ for four different reflector film thicknesses: (a) $W=10$ nm, (b) $W=20$ nm, (c) $W=40$ nm, and (d) $W=80$ nm.
@@
For silver CRs, the normalized loss factors displayed moderate deviations from the reference values except for very small values of $S/\lambda_{Ag}$, where a steep increase in the losses occurred due to stronger interactions with the nearby CR metal films when the thickness, $W$, was sufficiently large. In contrast, for gold CRs, Figure 10 indicates larger fluctuations in the loss factor around the reference levels.
@@
The calculated values of the Purcell factor, $f_{R}$, also exhibited a pronounced effect due to the interaction with the corner sheets when $W$ was sufficiently large. This interaction significantly impacted both silver and gold NAs, and the associated $S/\lambda_{M}$ range was larger than that observed for $f_{L}$ in silver NRs, increasing with $W$. The variation in $S/\lambda_{M}$ followed a distinct pattern: $f_{R}$ reached a maximum near the first peak of directivity and then dropped to a minimum around the directivity null at $S/\lambda_{M}=1$ before rising again near the second peak.
Figure 11 displays the radiation efficiency $\eta$ of the gold and silver CR NAs as a function of the feeder position $S$. As discussed in Section 2, this parameter is a very good approximation regarding the quantum efficiency of the NA for a highly efficient fluorophore with a negligible non-radiative transition rate. For thin metal sheets, with $W=10$ nm values, the fluctuations in $\eta$ as a function of $S$ were of small amplitudes. Larger efficiencies were observed for silver NAs, with values close to $\eta=0.6$, compared to gold NAs, which reached values around $\eta=0.4$. As the value of $W$ increased, the amplitude of the fluctuations in $\eta$ with $S$ became larger. Once again, the peak value of $\eta$ for silver, close to $0.7$, was higher than that for gold, which approached $0.6$.
@@
For silver CRs with $W=80$ nm values, we found that near the first directivity maximum at $S=0.30$ $\upmu$m ($S/\lambda_{Ag}=0.67$), the values were $D=20.1$, $f_{R}=14.5$, and $f_{L}=6.7$. The second maximum yielded a significantly higher directivity of $D=40.4$ at $S=0.55$ $\upmu$m ($S/\lambda_{Ag}=1.22$), with $f_{R}=6.0$ and $f_{L}=5.9$. Thus, the CR provided very large improvements in the directivity value when compared to those of the electric dipole emitter, with $D=1.5$.
@@
Figure 11: Radiation efficiency of silver (a) and gold (b) CR NA as a function of the normalized feeder position $S/\lambda_{M}$ for four different reflector film thicknesses $W$.
Considering the reference values for the NP emitter, $f_{R}=8.3$ and $f_{L}=5.4$, the first maximum offered a significant increase in both directivity and radiative emission rate enhancement factor, with only a modest rise in losses. At the second maximum, the directivity was much larger, though there was a slight reduction in $f_{R}$ in relation to the reference. The loss factor, however, remained close to that of the emitter without the CR. The gain figure of merit values were $G=13.8$ and $G=20.4$ for the first and second maxima, respectively, largely exceeding the value of $G=0.9$ obtained for the NA without the CR.
For the gold NA, the reference values were $f_{R}=4.9$ and $f_{L}=6.1$. With a $W=80$ nm thick. CR NA, at the first maximum, located at $S=0.35$ $\upmu$m ($S/\lambda_{Au}$=0.58), the directivity was $D=18.0$, with $f_{R}=8.2$ and $f_{L}=7.0$. The second maximum occurred at $S=0.75$ $\upmu$m ($S/\lambda_{Au}=0.58$), where $D=29.0$, $f_{R}=4.4$, and $f_{L}=6.7$.
For the gold CRs, the first maximum provided a high directivity and an improvement in the $f_{R}$ value. At the second maximum, while the directivity was even greater, the $f_{R}$ was close to the reference value. In both cases, the losses were similar to those of the gold NP without the CR. The gain figure of merit values were $G=8.0$ and $G=11.5$ for the first and second maxima, respectively, significantly exceeding the value of $G=0.7$ obtained for the NA without the CR but significantly in value lower than the values obtained for silver CR NAs. In addition to the impact of the better figures for the $f_{R}$ and $f_{L}$ values in the case of silver NAs, the lower values of $G$ obtained in the case of gold CR NAs were also due to the fact that $L/\lambda_{M}$ was smaller in the case of gold.
The incorporation of the CR in the proposed hybrid NAs is targeted to shape the radiation produced by the dipole-NP feeder system. It is important to evaluate the possible interaction effects between the CR and the feeder. First, we calculated the properties of a $W=80$ nm thick CR when the emitter-NP was replaced by a single dipole emitter feeder. Figure 12a,b compare the the directivities as a function of $S/\lambda_{M}$ obtained with the CR NA with and without the spherical NP coupled to the dipole emitter. In the latter case, the position of the dipole emitter was centered at the origin. The results demonstrate that the spatial shaping of the emission obtained by using the CR is largely independent of the dipole feeding system employed, since the differences between the directivities obtained with and without the NP were negligible. Figure 12c,d display the Purcell and loss factors for silver and gold nanoantennas, respectively, in the same cases as the previous figures. The losses due to the CRs were much smaller that those obtained for a dipole coupled to a metal nanosphere and showed little dependence with respect to the position of the dipole emitter along the corner axis. For the case of silver CRs, the losses were negligible at $\lambda=450$ nm. As regards the Purcell factor, we observed the same type of pattern as that shown in Figures 9 and 10 regarding the variation in $f_{R}$ with $S/\lambda_{M}$ due to the interaction of the dipole emitter with the CR plasmonic antenna, but the deviations from the unity were milder that those of the NP–emitter system and much smaller than those of the NP–emitter incorporated as a feeder system in the CR NA. In summary, whereas the CR alone was responsible for managing the directive properties of the hybrid NA, the strong plasmonic interactions observed in the system, with large variations in $f_{R}$, were found to be mediated by the spherical NP at the feeder arrangement.
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
Figure 12: Directivity of silver (a) and gold (b) CR antennas with and without NP coupled to the dipole emitter. Purcell ($f_{R}$) and loss ($f_{L}$) factors for silver (c) and gold (d) CRs with a single dipole emitter feeder.
Figure 13 displays the plasmonic resonance wavelengths when the NPs were incorporated into the CR setup. These results illustrate the impact of NP–CR plasmonic interactions on the Purcell factor. Note that the distances between the feeder NP and the corner apex are unnormalized in these plots. The results obtained for spherical NPs, shown in Figure 2, exhibit minimal variation in the resonance wavelengths with respect to particle size and the distance between the NP and the emitter. In contrast, Figure 13 demonstrates that, for sufficiently large mirror thicknesses $W$, there were significant shifts in the resonance wavelengths as the position of the feeder emitter–NP was moved along the corner axis. As $S$ increased, there was an initial rise in $\lambda_{res}$, followed by a sudden drop to approximately the resonance wavelength of the isolated NPs. This drop occurred for values of $S$ near the minimum $f_{R}$ in Figures 9 and 10, highlighting the relationship between these results and the interaction properties of the CR with the NP in close proximity.
The modification of the emitter’s fluorescence rate by the NA system is directly related to the local field enhancement (LFE) at the emitter’s position due to the surrounding plasmonic environment [21]. To evaluate the LFE factor, the NA setup must be in receiving mode. Reciprocity [15, 16] ensures that the antenna properties remain the same whether it is receiving or transmitting electromagnetic radiation. For these calculations, a linearly polarized plane wave, with the $E$ field aligned along the $Z$ axis, impinges on the NA along the $Y$ axis in the reverse direction. The local field enhancement factor can be evaluated as the ratio
$f_{E}=\dfrac{\left|\mathbf{E}\right|}{\left|\mathbf{E}_{0}\right|}$ (9)
between the field strength at the dipole source position $\left|\mathbf{E}\right|$ and the incident field strength $\left|\mathbf{E}_{0}\right|$, which is the field strength that would be found at that position in the absence of the NA.
Figure 14a displays the calculated electric field strength $|\mathbf{E}|$ over the $XZ$ plane scattered from a gold nanosphere when illuminated by a unit amplitude E-field plane wave. For the transmitting mode of the NA, the dipole source was positioned at $x=0$, $z=50$ nm. Figure 14b shows the corresponding electric field amplitude for the CR NA at the first directivity maximum. The results reveal a significant LFE for the isolated NP and a strong enhancement of this effect by the CR at the selected feeder position $S=0.35$ $\upmu$m in agreement with the values obtained for $f_{R}$. The calculated LFE factor for the previously optimized geometries yielded, for silver NAs, $f_{E}=13.8$ at $S=0.3$ $\upmu$m and $f_{E}=12.8$ at $S=0.55$ $\upmu$m. For gold NAs, the values were $f_{E}=9.9$ at $S=0.35$ $\upmu$m and $f_{E}=9.1$ at $S=0.75$ $\upmu$m. For the isolated NPs, the calculated LFE factors were $f_{E}=2.8$ for silver and $f_{E}=2.2$ for gold nanospheres.
(a) (b)
Refer to caption Refer to caption
Figure 13: Calculated resonance wavelengths of silver (a) and gold (b) NPs as a function of the absolute feeder position $S$ for four different reflector film thicknesses: $W=10$ nm, $W=20$ nm, $W=40$ nm, and $W=80$ nm.
(a) (b)
Refer to caption Refer to caption
Figure 14: Field strength $|\mathbf{E}|$ across the $XZ$ plane in the environment of a $d=70$ nm gold NP at the origin illuminated by an unit magnitude $Z$-polarized E-field plane wave. The plot (a) displays the results for the isolated NP and (b) when the NP was within the CR at $S=0.35$ $\upmu$m.
The properties of plasmonic NAs are largely influenced by interactions with nearby metal nanostructures. These can be leveraged in designing more efficient emitting structures, such as nanoparticle dimers [8, 9] or optimized NP arrays [13, 14], and are also relevant in the development of substrates for surface-enhanced fluorescence spectroscopy [39] and surface-enhanced Raman scattering (SERS) applications [40]. We now examine the potential effects on the emission properties of the CR NA due to a nearby metal NP. To this end, we considered the presence of a $d=70$ nm NP at a distance $\rho=1.5$ $\upmu$m from the feeder NP at the origin, which was positioned at various angles $\alpha$ from the $Y$ axis. The position vector of the NP is given by $\mathbf{r}_{n}=(\rho\sin(\alpha_{n}),\rho\cos(\alpha_{n}),0)$, where the values of $\alpha$ considered were $\alpha_{n}=n\cdot 30^{\circ}$ for $n=0\dots 6$.
Figure 15 shows the results calculated for a silver CR NA with $W=80$ nm in thickness and a feeder position of $S=0.55$ $\upmu$m in the presence of an interfering NP. The results indicate a very small deviation in the values of $f_{R}$ and $f_{L}$ compared to those determined without the interfering NP, as shown in Figure 9d, regardless of the position angle $\alpha$. The directivity values in Figure 15b also exhibit minimal deviations from those obtained without the interfering NP, shown in Figure 5a, especially when the interfering NP was positioned behind the CR. These calculations demonstrate a certain degree of robustness in the CR NA’s response to plasmonic NP contamination in the NA surroundings, at least outside the region defined by a circle centered at the feeder NP with a radius equal to the mirror length $L$ of the CR.
(a) (b)
Refer to caption Refer to caption
Figure 15: (a) Purcell, $f_{R}$, and loss, $f_{L}$, factors; (b) directivity of a $W=80$ nm, $L=1.5$ $\upmu$m CR silver NA with the feeder at $S=0.55$ $\upmu$m in the presence of an interfering $d=70$ nm silver NP at position $(\rho\sin(\alpha),\rho\cos(\alpha),0)$ for different values of $\alpha$.
@@
Figure 16 shows the directivities calculated for the silver and gold CR NAs when the corner length was increased to $L=3$ $\upmu$m for three values of $W$ ($10$ nm, $40$ nm, and $80$ nm). For this $L$, the peak directivities were essentially the same for the silver and gold CRs, even though the $L/\lambda_{M}$ value was larger for silver. This indicates a saturating effect regarding the increase in $D$ with $L$.
(a) (b)
Refer to caption Refer to caption
Figure 16: Values obtained of the directivity as a function of the feeder position $S$ normalized to the corresponding target wavelength for three different reflector film thicknesses when the length of the CR planes was increased to $L=3$ $\upmu$m. (a) Results for silver CR. (b) Results for gold CR.
@@
In the case of gold, the relative positions of the maxima and minima as a function of $S/\lambda_{M}$ closely matched those found for $L=1.5$ $\upmu$m, with a null occurring at $S=\lambda_{M}$ for $W=80$ nm. In contrast, for silver, significant shifts in the directivity maxima and minima as a function of $S/\lambda_{M}$ were observed , and these shifts were highly dependent on the value of $W$.
The normalized radiative emission and loss rates for $L=3$ $\upmu$m CRs are displayed in Figure 17 for the three values of $W$ considered. The results closely resemble those obtained for $L=1.5$ $\upmu$m, indicating that the impact of the CR on the loss and emission properties was due to the interactions occurring in a localized region near the apex of the CR.
Figure 18 shows the gain, $G$, of the silver and gold CR NAs as a function of $S/\lambda_{M}$. The gain provides a comprehensive measure of on-axis photon emission, incorporating all three key parameters: $f_{R}$, $f_{L}$, and $D$. For the gold CR NAs, the maximum gain occurred at a metal thickness of $W=80$ nm. The first gain peak, with a value of $G=11.8$, was found at $S/\lambda_{Au}=0.67$ and $S/\lambda_{Au}=0.75$. The second peak, $G=21.8$, was located at $S/\lambda_{M}=1.33$.
@@
Figure 17: Radiative transition rate enhancement $f_{R}$ and normalized loss $f_{L}$ factors of a gold CR with $L=3$ $\upmu$m as a function of the feeder position $S/\lambda_{M}$ for three different reflector film thicknesses: (a) silver CR $W=10$ nm, (b) gold CR $W=10$ nm, (c) silver CR $W=40$ nm, (d) gold CR $W=40$ nm, (e) silver CR $W=80$ nm, (f) gold CR $W=80$ nm.
@@
Figure 18: Nanoantenna gain as a function of the feeder position $S$ normalized to the corresponding target wavelength for three different reflector film thicknesses when the length of the CR planes was increased to $L=3$ $\upmu$m. (a) Results for silver CR. (b) Results for gold CR.
In the case of the silver CRs, the aforementioned shift of the directivity maxima with respect to $S/\lambda_{Ag}$ resulted in the highest directivity for $W=40$ nm rather than for $W=80$ nm. This shift aligns the directivity peak with a higher $f_{R}$ value at $W=40$ nm, leading to the maximum gain being achieved at this thickness. The first gain peak for silver was $G=12.3$ at $S/\lambda_{Ag}=0.5$, and the second was $G=30.0$ at $S/\lambda_{Ag}=1.25$.
The directivity values obtained for the CR NA were, in all cases, well below the bounds set by the Kildal limit for the CR sizes considered [28]. Nevertheless, the proposed scheme offers a significant improvement over a single NP emitter while maintaining high implementation simplicity, which is consistent with the advantages of CR antennas used in radio frequency applications. Although superdirective dielectric NAs would allow for a more compact setup, the proposed scheme enables the full exploitation of plasmonic nanostructures’ potential [34].
@@ 4 Conclusions
The analysis of dipole–NP interactions, combined with CR integration, reveals important insights into the enhancement of radiative properties for silver and gold NPs. Knowing the fundamental behavior of NP interactions is crucial for understanding the modifications introduced by the incorporation of the CR in the system. Initially, isolated dipole–NP calculations without the CR showed trends consistent with previous theoretical and experimental studies, particularly regarding the variation in resonance wavelength, the radiative emission rate enhancement factor, and loss factor with silver and gold particle size and dipole emitter separation. This preliminary study established a set of parameters for the subsequent analysis of the CR. The results also suggest that silver NPs are more effective than gold NPs in enhancing radiative emission while maintaining lower loss factors.
Integrating CRs into silver and gold NA systems leads to substantial improvements in directivity. The variation in directivity with respect to the normalized distance $S/\lambda_{M}$ reveals that thicker metal films generally lead to more regular oscillatory patterns in directivity. At the first directivity maxima, significant gains in the radiative enhancement factor, $f_{R}$, were also achieved with minimal increases in the loss factor, $f_{L}$, leading to high gain figures of merit, $G$. At the second directivity maximum, the CR mainly enhanced directivity, with the highest values of $D$ corresponding to the global maxima of $G$.
Extending the corner length to $L=3$ $\upmu$m demonstrated a saturation effect in the enhancement of directivity for both materials, with similar peak values achieved for the gold and silver CRs despite differences in their $L/\lambda_{M}$ values. Meanwhile, the patterns for the variations in $f_{R}$ and $f_{L}$ closely resembled those found at $L=1.5$ $\upmu$m. These findings highlight the localized nature of CR-induced enhancements near the apex. However, silver CRs exhibited larger shifts in their positions of the directivity maxima and minima with changes in $W$. This resulted in the maximum gain for silver CR NAs at a metal thickness of $W=40$ nm, rather than $W=80$ nm, due to the alignment of the directivity maximum with a higher value of $f_{R}$.
Overall, the results indicate that silver CR NAs outperform gold CR NAs in terms of radiative enhancement and gain, offering more efficient light emission and directivity control for nanoscale optical applications, although different spectral windows are targeted by the choice of NA metal.
The reflector angle selected for the analysis aligns with a typical value used in RF applications, corresponding to optimization for operation at the more favorable second gain maximum. The analysis reveals a strong metal–feeder interaction near the corner apex, suggesting the potential benefits of further exploring corner geometries to optimize the first gain maximum, which will be addressed in future works.